# Multi-perspective Feedback-attention Coupling Model for Continuous-time Dynamic Graphs

### v1

@@ V-B Baselines
We select nine temporal models (JODIE [24], DyRep [21], TGAT [26], TGN [25], CAWN[29], EdgeBank[30], TCL[27], GraphMixer[39], and DyGFormer[28]) as benchmarks. Additionally, we enhance two superior static graph models, GAT and GraphSAGE, to temporal models (GAT-t and GraphSAGE-t) and use them as new baselines. Refer to the Related Work subsection for baseline details. To ensure fair comparisons, all baselines share consistent settings across eight datasets: batch size (200), number of temporal neighbors (10), embedding dimension (172), and dropout rate chosen from (0.0, 0.1, 0.2, 0.3, 0.4). The learning rate is set to 0.0001. The remaining baseline settings align with those specified in their respective papers. The extension of GAT and GraphSAGE to temporal models involves incorporating temporal information.
@@ V-C Experimental setup and evaluation metrics
For fair comparisons, our model and baselines share identical hyperparameters whenever possible. Across all datasets and experiments, we maintain a batch size of 200, a learning rate of 0.0001, and an embedding dimension of 172. In the link prediction task, we run experiments on six temporal datasets. During training, we compute event probabilities using an equal number of negative and positive samples. Reference metrics include average precision (AP), accuracy (ACC), and area under the ROC curve (AUC). For the node classification task on the benchmark datasets (Wikipedia and Reddit), AUC is employed due to label imbalances. Temporal neighbors, selected from the 10 nearest neighbors in chronological order, are used in all experiments. The final results are the average of 10 model runs. For all tasks, we split each dataset chronologically into training (75%), validation (15%), and test (15%) sets.
V-D Experiments
@@ V-D1 Dynamic link prediction
We conduct experiments on six temporal datasets, i.e. USLegis, ComplexDynamics, SocialEvo, Mooc, UNVote, LastFM for dynamic link prediction. The task is to determine, given nodes $v_{i}$ and $v_{j}$ at time $t$, whether they will interact with each other? Within this task, we design two additional subtasks, the inductive and the transductive subtasks. The transductive task indicates that the nodes appearing in the test set were seen by the model during training, while they were not seen in the inductive task.
From the results of dynamic link prediction on the six datasets (see Tables I, II, III, and IV), our proposed algorithm demonstrates a distinct advantage in this task. Across four datasets—MOOC, UNvote, USLegis, and LastFM—MPFA consistently outperforms other algorithms in both transductive and inductive subtasks based on evaluation criteria such as AP, AUC, and ACC. For instance, in the transductive task on the MOOC dataset (see Table I), MPFA achieves an AP score of 92.98, surpassing the second-best algorithm (GAT-t) by 3.61 percentage points and outperforming classical algorithms TGAT and TGN by 7.9 and 4.09 percentage points, respectively. MPFA also exhibits advantages in SocialEvo and ComplexDynamics, particularly in the AP scoring criterion (refer to Table II). However, on SocialEvo and ComplexDynamics, MPFA slightly lags in ACC and AUC. This discrepancy may be attributed to a frequency dominance feature in these datasets, favoring algorithms like DyGFormer and TCL. Examining the two subtasks, both MPFA and the baselines demonstrate a stronger advantage in the transductive task across most datasets. This is primarily due to the fact that, in the transductive task, the model predicts nodes seen during training, whereas the inductive task involves predicting unseen nodes. An interesting anomaly is observed in the LastFM dataset, where the inductive task outperforms the transductive task. Taking MPFA as an example, the AP, AUC, and ACC for the inductive task are 88.05, 88.55, and 81.60, respectively, while the corresponding values for the transductive task are only 84.23, 84.78, and 76.99 (refer to Table II and IV). In summary, our proposed continuous-time dynamic graph modeling algorithm proves to be effective and efficient in predicting dynamic links.
@@
Figure 4: Ablation experiment results. Results of MOOC ablation experiments are denoted by (a) and (b), and for the USLegis dataset by (c) and (d). W/O RP: without Raw Perspective; W/O EP: without Evolving Perspective; W/O RED: Raw and Evolving perspectives without Dynamic update; W/O ED: Evolving perspective without Dynamic update.
@@
Figure 5: Dynamic link prediction results for six models with varying numbers of history neighbors on MOOC and USLegis reveal that superior predictions with fewer history neighbors suggest the model requires only a limited number of interactions for effective long-term dependency properties.
@@ V-D2 Dynamic node classification
We perform the dynamic classification experiment on two datasets with dynamic labels (Wikipedia and Reddit). The purpose of this experiment is to determine whether a node is banned after going through an event. For example, on Wikipedia, when a user posts, the platform determines whether to ban them based on their posts. After obtaining the temporal network embeddings, we train an MLP decoder to predict node label changes. Due to the label imbalance in the dataset, we only use AUC as the evaluation metric in this experiment. For dynamic node classification, Table V compares the results of MPFA and baselines, where our method achieves state-of-the-art results on both benchmark datasets. We outperform the second best method by 2.9 and 4.3 on the Wikipedia and Reddit datasets, respectively. Thus, our proposed model is also valid for the dynamic classification task.
@@ V-D3 Ablation studies
To evaluate the effectiveness of different components of our proposed model, we conduct the ablation studies in this part. We experiment with different combinations of modules from MPFA for dynamic link prediction on the MOOC and USLegis datasets, aiming to observe their specific properties (see Figure 4 for details). From the results of the overall ablation experiments on both datasets, MPFA with two perspectives achieves the highest performance in both transductive and inductive tasks, as evidenced by superior values in AP, AUC, and ACC. This result substantiates the effectiveness of MPFA. Conversely, the evolving perspective without dynamic updating emerges as the least effective, as indicated by the lowest values across all three evaluation criteria. This underscores the significance of the dynamic updating module designed to capture long-term dependencies. Regarding the evolving and raw perspectives, the model exclusively featuring the evolving perspective demonstrates the second-best performance on MOOC, while on USLegis, the model with only the raw perspective achieves the second-best performance. This discrepancy highlights the varying roles of these perspectives in different scenarios. Furthermore, the model incorporating both evolving and raw perspectives but lacking dynamic updating performs suboptimally, underscoring the crucial role of long-term dependency in effective model learning.
@@ V-D4 Long dependency propertity
To verify that our model can maintain the long-term dependency property with a smaller number of historical neighbors, we conduct experiments on link prediction performance under different numbers of temporal neighbors for MPFA and five baselines on the MOOC and USLegis datasets. As shown in Figure 5, we observe that MPFA maintains a stable state when the number of temporal neighbors varies from 1 to 30, both on the MOOC and USLegis datasets. In comparison to other models, especially on the MOOC dataset, all five baselines exhibit the characteristic of performance improvement with an increasing number of neighbors. Among them, CAWN shows the most significant performance improvement. Although these baselines occasionally display jumps in predicted values for the transductive and inductive subtasks of USLegis, they still require an increase in the number of neighbors to achieve performance improvement. In summary, our proposed dynamic update module is effective. Superior predictions with fewer history neighbors suggest that our model requires only a limited number of interactions for effective long-term dependency properties.
@@ V-D5 Hyperparametric analysis
In this section, we examine the performance of MPFA under three main hyperparameters on two datasets, MOOC and USLegis. As illustrated in Figure 6, MPFA’s performance remains relatively stable with variations in batch size, embedding dimension, and the number of temporal neighbors. In Figure 6(a), the model performance exhibits a slight decrease as the batch size increases. In Figure 6(b), there is a slight increasing trend in model performance with higher embedding dimensions. Figure 6(c) shows the model remains stable up to 30, with a slight decrease thereafter. In USLegis, MPFA experiences two significant fluctuation points with different batch sizes: 200 points in the transductive task and 300 points in the inductive task (Figure 7(a)). Across varying embedding dimensions (Figure 7(b)), similar to the MOOC dataset, MPFA’s performance tends to increase slightly. In experiments with different numbers of temporal neighbors (Figure 7(c)), performance in the transductive task remains stable, with some fluctuations observed in the inductive task.
@@ V-D6 Attention score visualization
In this section, we visualize attention coefficients for both perspectives on the MOOC and USLegis datasets (see Figure 8). This enables observation of the focus areas for the two perspectives in transductive and inductive tasks. Generally, the evolving perspective exhibits higher attention coefficients, typically reaching values of 0.3 or higher, while the raw perspective tends to have smaller coefficients, usually below 0.4. These coefficient values align with the model’s ability to gain deeper insights into historical events post-learning, emphasizing extraction of a limited amount of valid information. On both the MOOC and USLegis datasets, the evolving perspective in the transductive task directs more attention to neighbors closer in the time series. Specifically, on the MOOC dataset, the evolving perspective exhibits exceptional attention to the nearest temporal neighbors. Conversely, in the inductive task, the focus is relatively more diffuse. For the raw perspective on both datasets, temporal attention on the MOOC dataset is notably larger and broader in the transductive task. On the USLegis dataset, attention remains essentially the same for both transductive and inductive tasks.
@@ V-D7 Time efficiency
To assess the runtime efficiency of MPFA and current leading baselines, we select five algorithms (TCL, TGAT, GraphMixer, CAWN, DyGFormer) to evaluate the time consumption for achieving optimal training results on the MOOC and USLegis datasets, as depicted in Figure 9. Although TGAT requires the least time to optimize its model parameters on MOOC, it does not perform well in terms of link prediction results. In contrast, MPFA, despite consuming more time than TGAT, achieves optimal test performance on MOOC. Similarly, TCL also has the shortest training time but fails to yield better test results. We attribute this to both models potentially inadequately fitting the dataset, causing premature training termination. Overall, MPFA demonstrates an advantageous balance between runtime efficiency and optimal performance.
V-E Discussion
Graph neural networks have proven highly effective for processing graph-structured data, with numerous algorithms proposed for various scenarios of static graphs. However, in real-world applications, network structures are not static but constantly evolve over time. Consequently, conventional static graph expression algorithms struggle to adapt to these dynamically changing scenarios. Several models have been introduced to address dynamic graph problems, categorized into discrete and continuous models. Discrete models, which handle network evolution through snapshots extracted at specific time intervals, suffer from information loss due to their event stream representation. Continuous dynamic graph models, while capturing the moment of evolution, often neglect the raw interaction state, diminishing the impact of historical events on predicting future events. Moreover, many current continuous-time dynamical graph models excessively consume time resources to obtain long-term dependencies. To overcome these challenges, this paper proposes distinct treatment schemes:
Handling Changing Topology: For static graphs unable to adapt to changing topology, we transform the dynamic graph network into an event stream, creating a fine-grained model for comprehensive consideration of network evolution details. Simultaneously, we assess whether classical static models maintain an advantage when extended to temporal graphs by adapting GAT and GraphSAGE. Despite performance differences compared to MPFA, experimental results show they outperform many dynamic models.
Incorporating Multiple Perspectives: To address the limitation of existing continuous-time dynamic models focusing solely on the evolving perspective, we propose incorporating a raw perspective to observe the network’s evolution from multiple directions. Ablation experiment results highlight distinct roles for each perspective in various scenarios. By coupling the evolving and raw perspectives, we enhance the model’s applicability to different contexts.
Dynamic Updating Model for Long-Term Dependencies: Existing models typically achieve long-term dependencies by increasing the number of historical neighbors. In contrast, we adopt a dynamic updating model, allowing the use of a small number of historical neighbors to acquire long-term dependencies.
Despite achieving good results, our model faces certain challenges. For instance, when handling networks with a small number of nodes (a few hundred nodes) and millions of edges, its performance tends to be slightly lower than some contemporary models. Additionally, increasing the batch size to a large value results in a decrease in model performance. These issues need to be addressed in future work.
@@ VI Conclusion
In this paper, we address dynamically changing graph networks through event streams, proposing an effective continuous-time dynamic graph model. Our model pioneers the exploration of dynamic network evolution from both the raw and evolving perspectives. The evolving perspective captures real-time network dynamics, revealing fine-grained topology transformations and long-term dependencies with fewer historical interactions. Ablation experiments and long-term dependency characterization confirm the effectiveness of the evolving perspective. The often-overlooked raw perspective, representing the nature of interaction events, shows its advantages in diverse datasets or domains through ablation experiments. Relying on only one perspective may weaken the model’s generalization ability, emphasizing the importance of simultaneous learning from both perspectives. Validation through dynamic link prediction, dynamic classification, and numerous experiments underscores the effectiveness of our proposed model. This approach introduces a novel idea for future research and holds promise for application across multiple scenarios with dynamic changes.

### v2

@@ V-B Baselines
We select nine temporal models (JODIE [24], DyRep [21], TGAT [26], TGN [25], CAWN[29], EdgeBank[30], TCL[27], GraphMixer[39], and DyGFormer[28]) as benchmarks. Additionally, we enhance two superior static graph models, GAT and GraphSAGE, to temporal models (GAT[8]-t and GraphSAGE[9]-t) and use them as new baselines. Refer to the Related Work subsection for baseline details. To ensure fair comparisons, all baselines share consistent settings across eight datasets: batch size (200), number of temporal neighbors (10), embedding dimension (172), and dropout rate chosen from (0.0, 0.1, 0.2, 0.3, 0.4). The learning rate is set to 0.0001. The remaining baseline settings align with those specified in their respective papers. The extension of GAT and GraphSAGE to temporal models involves incorporating temporal information.
@@ V-C Experimental setup and evaluation metrics
For fair comparisons, our model and baselines share identical hyperparameters whenever possible. Across all datasets and experiments, we maintain a batch size of 200, a learning rate of 0.0001, and an embedding dimension of 172. During training, we compute event probabilities using an equal number of negative and positive samples. Reference metrics include average precision (AP), accuracy (ACC), and area under the ROC curve (AUC). For the node classification task on the benchmark datasets (Wikipedia and Reddit), AUC is employed due to label imbalances. Temporal neighbors, selected from the 10 nearest neighbors in chronological order, are used in all experiments. The final results are the average of 10 model runs. For all tasks, We perform a chronological split, assigning 70% for training, 15% for validation, and 15% for testing, based on node interaction timestamps. All experiments are conducted under the Ubuntu 18.04 system, the Pytorch deep learning framework, and the NVIDIA TITAN XP GPU 12GB.
V-D Experimental results and analysis
@@ V-D1 Dynamic link prediction
We perform the dynamic link prediction task on six dynamic datasets (USLegis, ComplexDynamics, SocialEvo, Mooc, UNVote, LastFM). The task is to determine whether given nodes $v_{i}$ and $v_{j}$ will interact with each other at time $t$? In this task, we design two additional subtasks: an inductive subtask and a transductive subtask. The transductive subtask means that the nodes present in the test set were seen by the model during training, but were not seen in the inductive subtask.
From the results of dynamic link prediction on the six datasets (see Tables I, II, III, and IV), our proposed algorithm demonstrates a distinct advantage in this task. Across four datasets—MOOC, UNvote, USLegis, and LastFM—MPFA consistently outperforms other algorithms in both transductive and inductive subtasks. For instance, in the transductive task on the MOOC dataset (see Table I), MPFA achieves an AP score of 92.98%, surpassing the second-best algorithm (GAT-t) by 3.61% and outperforming classical algorithms TGAT and TGN by 7.9% and 4.09%, respectively. MPFA also exhibits advantages in SocialEvo and ComplexDynamics, particularly in the AP scoring criterion (refer to Table II). However, on SocialEvo and ComplexDynamics, MPFA slightly lags in ACC and AUC. This discrepancy may be attributed to a frequency dominance feature in these datasets, favoring algorithms like DyGFormer and TCL. Examining the two subtasks, both MPFA and the baselines demonstrate a stronger advantage in the transductive task across most datasets. This is primarily due to the fact that, in the transductive task, the model predicts nodes seen during training, whereas the inductive task involves predicting unseen nodes. An interesting anomaly is observed in the LastFM dataset, where the inductive task outperforms the transductive task. Taking MPFA as an example, the AP, AUC, and ACC for the inductive task are 88.05%, 88.55%, and 81.60%, respectively, while the corresponding values for the transductive task are only 84.23%, 84.78%, and 76.99% (refer to Table II and IV). In summary, our proposed continuous-time dynamic graph modeling algorithm proves to be effective and efficient in predicting dynamic links.
@@
Figure 4: Results of ablation experiments. The results of the MOOC ablation experiments are denoted by (a) and (b), and for the USLegis dataset by (c) and (d). W/O RP: without Raw Perspective; W/O EP: without Evolving Perspective; W/O RED: Raw and Evolving perspectives without Dynamic update; W/O ED: Evolving perspective without Dynamic update.
@@
Figure 5: Long-term dependency effects of six models on MOOC and USLeigs datasets under different numbers of neighbors.
@@ V-D2 Dynamic node classification
We perform the dynamic classification experiment on two datasets with dynamic labels (Wikipedia and Reddit). The purpose of this experiment is to determine whether a node is banned after going through an event. For example, on Wikipedia, when a user posts, the platform determines whether to ban them based on their posts. After obtaining the temporal network embeddings, we train an MLP decoder to predict node label changes. Due to the label imbalance in the dataset, we only use AUC as the evaluation metric in this experiment. For dynamic node classification, Table V compares the results of MPFA and baselines, where our method achieves state-of-the-art results on both benchmark datasets. We outperform the second-best method by 2.9% and 4.3% on the Wikipedia and Reddit datasets, respectively. Thus, our proposed model is also valid for the dynamic classification task.
@@ V-D3 Ablation studies
To evaluate the effectiveness of different components of our proposed model, we conduct the ablation studies in this part. We experiment with different combinations of modules from MPFA for dynamic link prediction on the MOOC and USLegis datasets, aiming to observe their specific properties (see Figure 4 for details). From the results of the overall ablation experiments on both datasets, MPFA with two perspectives achieves the highest performance in both transductive and inductive tasks, as evidenced by superior values in AP, AUC, and ACC. This result substantiates the effectiveness of MPFA. Conversely, the evolving perspective without dynamic updating emerges as the least effective, as indicated by the lowest values across all three evaluation criteria. This underscores the significance of the dynamic updating module designed to capture long-term dependencies. Regarding the evolving and original perspectives, the model exclusively featuring the evolving perspective demonstrates the second-best performance on MOOC, while on USLegis, the model with only the raw perspective achieves the second-best performance. This discrepancy highlights the varying roles of these perspectives in different scenarios. Furthermore, the model incorporating both evolving and original perspectives but lacking dynamic updating performs suboptimally, underscoring the crucial role of long-term dependency in effective model learning.
@@ V-D4 Long dependency propertity
To verify that our model can maintain the long-term dependency property with a smaller number of historical neighbors, we perform experiments on the link prediction performance under different numbers of temporal neighbors for MPFA and five baselines on the MOOC and USLegis datasets. As shown in Figure 5, we observe that MPFA maintains a stable state when the number of temporal neighbors varies from 1 to 30, both on the MOOC and USLegis datasets. Compared to other models, especially on the MOOC dataset, all five baselines show the characteristic of performance improvement with increasing number of neighbors. Among them, CAWN shows the most significant performance improvement. Although these baselines occasionally show jumps in the predicted values for the transductive and inductive subtasks of USLegis, they still require an increase in the number of neighbors to achieve performance improvement. In summary, our proposed dynamic update module is effective. Superior predictions with fewer history neighbors suggest that our model requires only a limited number of interactions for effective long-term dependency properties.
@@ V-D5 Hyperparametric analysis
In this section, we examine the performance of MPFA under three main hyperparameters on two datasets, MOOC and USLegis. As shown in Figure 6, the performance of MPFA remains relatively stable with variations in batch size, embedding dimension, and number of temporal neighbors. In Figure 6(a), the model performance shows a slight decrease as the batch size increases. In Figure 6(b), there is a slightly increasing trend in model performance with higher embedding dimensions. Figure 6(c) shows that the model remains stable up to 30, with a slight decrease thereafter. In USLegis, MPFA experiences two significant fluctuation points with different batch sizes: 200 points in the transductive task and 300 points in the inductive task (Figure 7(a)). With different embedding dimensions (Figure 7(b)), similar to the MOOC dataset, the performance of MPFA tends to increase slightly. In experiments with different numbers of temporal neighbors (Figure 7(c)), performance remains stable in the transductive task, while some fluctuations are observed in the inductive task.
@@ V-D6 Attention score visualization
In this section, we visualize the attention coefficients for both perspectives on the MOOC and USLegis datasets (see Figure 8). This allows us to observe the areas of focus for the two perspectives in transductive and inductive tasks. In general, the evolving perspective has higher attention coefficients, typically reaching values of 0.3 or higher, while the original perspective tends to have smaller coefficients, typically below 0.4. These coefficient values are consistent with the model’s ability to gain deeper insights into historical events after learning, emphasizing the extraction of a limited amount of valid information. On both the MOOC and USLegis datasets, the evolving perspective in the transductive task directs more attention to neighbors closer in the time series. Specifically, on the MOOC dataset, the evolving perspective shows exceptional attention to the closest temporal neighbors. Conversely, in the inductive task, the focus is relatively more diffuse. For the original perspective on both datasets, temporal attention is significantly larger and broader on the MOOC dataset in the transductive task. For the USLegis dataset, attention remains essentially the same for both transductive and inductive tasks.
@@ V-D7 Time efficiency
To evaluate the runtime efficiency of MPFA and the current leading baselines, we select five algorithms (TCL, TGAT, GraphMixer, CAWN, DyGFormer) to evaluate the time taken to achieve optimal training results on the MOOC and USLegis datasets, as shown in Figure 9. Although TGAT requires the least time to optimize its model parameters on MOOC, it does not perform well in terms of link prediction results. In contrast, although MPFA takes more time than TGAT, it achieves optimal test performance on MOOC. Similarly, TCL has the shortest training time, but does not produce better test results. We attribute this to the fact that both models may have an inadequate fit to the dataset, leading to premature training termination. Overall, MPFA demonstrates a favorable balance between runtime efficiency and optimal performance.
@@ VI Conclusion
In this paper, we address dynamically changing graph networks through event streams, proposing an effective continuous-time dynamic graph model. Our model pioneers the exploration of dynamic network evolution from both the original and evolving perspectives. The evolving perspective captures real-time network dynamics, revealing fine-grained topology transformations and long-term dependencies with fewer historical interactions. Ablation experiments and long-term dependency characterization confirm the effectiveness of the evolving perspective. The often-overlooked original perspective, representing the nature of interaction events, shows its advantages in diverse datasets or domains through ablation experiments. Relying on only one perspective may weaken the model’s generalization ability, emphasizing the importance of simultaneous learning from both perspectives. Validation through dynamic link prediction, dynamic classification, and numerous experiments underscores the effectiveness of our proposed model. This approach introduces a novel idea for future research and holds promise for application across multiple scenarios with dynamic changes. In the future, we will explore extending the model to larger data sets and require other forms of feedback attention.